# Functoriality in Finitary Vacuum Einstein Gravity and Free Yang-Mills

### 24 Jan 2024 v1

@@ 2.3.1 Interregnum: Drawing Formal Links with 3rd Quantisation
Could the generation/emergence of (inertio-gravitational) mass be somehow accounted for by a (spontaneous) symmetry breaking-type of mechanism, whereby, the dynamical automorphism group ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ of the ADG-gravitational field $(\mathbf{\mathcal{E}},\mathcal{D})$ reduces to its subroup that leaves $\mathrm{ker}(\mathcal{U})$ invariant? Alternatively intuited, could the emergence of inertio-gravitational mass be thought of as the result of some kind of ‘quantum anomaly’ of 3rd-quantized Vacuum Einstein Gravity? [^98] The epithet ‘quantum’ adjoined to ‘anomaly’ is intended to distinguish the effect intuited above from the usual anomalies. A ‘quantum anomaly’ is the ‘converse’ of an anomaly in the usual sense, in that what was a symmetry of the quantum theory (:an element of ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ in our case) ceases to be a symmetry of the ‘classical domain’ of our theory (:$\mathrm{ker}(\mathcal{U})$). Let it be stressed that the emergence of gravito-inertial ‘mass’ in the sense intuited here has a truly relational (:algebraic) and ‘global’ flavour reminiscent of Mach’s ideas: ‘global’ gravitational field symmetries in ${\mathcal{A}}ut\mathbf{\mathcal{E}}$ are locally reduced to inertial ones, and sheaf theory’s ability to interplay between local and global comes in handy in this respect [61]. (In [61], Mallios and Zafiris do a great job in highlighting exactly how sheaf theory allows one to go from ‘local’ to ‘global’, and vice versa.)

### 27 Jan 2024 v2

@@ 2.3.1 Interregnum: Drawing Formal Links with 3rd Quantisation
Could the generation/emergence of (inertio-gravitational) mass be somehow accounted for by a (spontaneous) symmetry breaking-type of mechanism, whereby, the dynamical automorphism group ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ of the ADG-gravitational field $(\mathbf{\mathcal{E}},\mathcal{D})$ reduces to its subroup that leaves $\mathrm{ker}(\mathcal{U})$ invariant? Alternatively intuited, could the emergence of inertio-gravitational mass be thought of as the result of some kind of ‘quantum anomaly’ of 3rd-quantized Vacuum Einstein Gravity? [^98] The epithet ‘quantum’ adjoined to ‘anomaly’ is intended to distinguish the effect intuited above from the usual anomalies. A ‘quantum anomaly’ is the ‘converse’ of an anomaly in the usual sense, in that what was a symmetry of the quantum theory (:an element of ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ in our case) ceases to be a symmetry of the ‘classical domain’ of our theory (:$\mathrm{ker}(\mathcal{U})$). Let it be stressed that the emergence of gravito-inertial ‘mass’ in the sense intuited here has a truly relational (:algebraic) and ‘global’ flavour reminiscent of Mach’s ideas: ‘global’ gravitational field symmetries in ${\mathcal{A}}ut_{\mathbf{A}}\mathbf{\mathcal{E}}$ are locally reduced to inertial ones, and sheaf theory’s ability to interplay between local and global comes in handy in this respect [61]. (In [61], Mallios and Zafiris do a great job in highlighting exactly how sheaf theory allows one to go from ‘local’ to ‘global’, and vice versa.)